Circular No. 6772

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

X1354–644 R. Remillard, F. Marshall, and T. Takeshima report for the RXTE All

Sky Monitor (ASM) Team at Massachusetts Institute of Technology and Goddard Space Flight Center: “The Rossi X-ray Timing Explorer’s ASM has detected a modest outburst from the recurrent transient and blackholecandidate binary, X1354–644. Renewed x-ray activity began around Nov. 1, and the average flux over the interval Nov. 1–9 was 16 mCrab at 2– 12 keV, increasing to 26 mCrab during Nov. 11–16 and to 50 mCrab on Nov. 17–18. An RXTE PCA observation between Nov. 18.02 and 18.11 UT found an average flux of 50 mCrab (2–10 keV). There were aperiodic variations, with a range of time scales from subseconds to 1 min and flare amplitudes as large as a factor of 3. We encourage radio, optical, and infrared observations of the counterpart, BW Cir (cf. IAUC 4357).”

CATACLYSMIC VARIABLE IN PEGASUS Y.-l. Qiu, Q.-y Qiao, and J.-y. Hu, Beijing Astronomical Observatory

(BAO); and A. Esamdin, Urumqi Astronomical Station, in the course of the BAO supernova survey with the 0.6-m telescope at Xinglong Station, have observed another outburst of the cataclysmic variable star reported on IAUC 6746. On Nov. 14.44 UT, the star was at R = 15.0; it was not present on CCD images obtained on Nov. 12 (limiting mag ∼ 17.5).

LEONID METEORS 1997 Reports of moderate Leonid counts have been received from many vi-

sual observers and suggest that activity peaked around Nov. 17.5–17.6 UT. Leonid fireballs of mag −4 to −9 were numerous around this time. P. Brown, observing from Edwards Air Force Base, CA, and D. H. Levy, Tucson, AZ, observed a peak of ∼ 1–2 Leonids/min around Nov. 17.505, after which time the rate dropped despite darker skies and higher radiant. S. J. O’Meara, Volcano, HI, reported seeing several bursts (3–5 in a few seconds) of bright Leonids visually at Nov. 17.56 and 17.63, and he observed an overall peak (∼ 40/hr) around Nov. 17.52–17.57; about a fourth of the Leonids were brighter than mag 1. E. P. Bus, Groningen, The Netherlands, reports that monitoring of VHF radio waves during Nov. 17.2–17.5 yielded many long-duration meteor echoes (> 20 s). K. Suzuki, Toyokawa, Aichi, Japan, reports that 50-MHz HAM radio signals were monitored continuously from Nov. 14.5 to 18.0, yielding a peak rate of ∼ 60–80 meteors/hr during Nov. 17.7–17.9, hinting at a possible second activity peak.

1997 November 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green